OpenStudy (anonymous):

What is 8 1/5 - 4 2/5 equal please help

OpenStudy (anonymous):

1/5 is .20 amd 2/5 is .4

OpenStudy (anonymous):

so use those instead of the fractions to make it easier

OpenStudy (anonymous):

No it doesn't make them into improper fraction

OpenStudy (anonymous):

Do you know ?? \[8\frac{1}{5} = ??\]

OpenStudy (anonymous):

41/5

OpenStudy (anonymous):

and do the same for the other

OpenStudy (anonymous):

so its 22/5

OpenStudy (anonymous):

Then whatn

OpenStudy (anonymous):

Now: \[\frac{41}{5} - \frac{22}{5}\]

hero (hero):

8 1/5 - 4 2/5 (8 - 4) + (1/5 - 2/5) (7 - 4) + (6/5 - 2/5) 3 + 4/5

OpenStudy (anonymous):

equals 19

hero (hero):

3 + 4/5 = 19/5

OpenStudy (anonymous):

19 is in the numerator..

hero (hero):

3 + 4/5 means 3 "and" 4/5

OpenStudy (anonymous):

then what what is the answer is it just 9

OpenStudy (anonymous):

i mean 19

hero (hero):

face palm...."shakes head"

OpenStudy (anonymous):

the waterineyes girl is eaiser to understand

OpenStudy (anonymous):

wait is it 19/5

OpenStudy (anonymous):

Someone help me

OpenStudy (anonymous):

Are you sure it is 19/5??

OpenStudy (anonymous):

And lastly, I am not a Girl..

hero (hero):

Are you sure it isn't 19/5?

OpenStudy (anonymous):

no it 3 4/5 right

hero (hero):

3 4/5 = 19/5 I already said that once.

hero (hero):

@snowballhockey, do you know what the equal sign means?

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

and yeah so that is the correct answer

OpenStudy (anonymous):

can someone tell me the right answer

OpenStudy (anonymous):

@hero

hero (hero):

The correct answer has been stated several times. I am sorry you do not understand well enough to realize that.

OpenStudy (anonymous):

ok well i gave u the best responce